at a certain temperature, the following reaction has an equilibrium constant of Kp=401

At a certain temperature, the following reaction h
Answer #1

PCl3 + Cl2 -----> PCl5

Intial Pressure 0 0 0.078 bar

Change P -x -x +x bar

Equli P -x -x 0.078 + x bar

Kp = (0.0780 + x) / x2 = 401

401x2 -x -0.078 = 0

x = 0.015249

So total pressure at equilibrium =  0.015249 (PCl3) +  0.015249 (Cl2) +  [0.015249+0.078] (PCl5)

total pressure at equilibrium = 0.1237 bar